# What Matters More For In-Context Learning under Matched Compute Budgets: Pretraining on Natural Text or Incorporating Targeted Synthetic Examples?

### v1

@@ Abstract
Does explicitly exercising the induction circuit during pretraining improve in-context learning (ICL), or is natural text sufficient, when compute is held constant (iso-FLOPs)? To test whether targeted synthetic data can accelerate the emergence of induction heads and enhance ICL performance, we introduce Bi-Induct, a lightweight curriculum that injects forward-copy (Induction), backward-copy (Anti, as a control), or a balanced mix, into the pretraining stream. We conduct iso-FLOPs pretraining across models from 0.13B to 1B parameters, evaluating effects across three axes: (i) few-shot performance on ICL benchmarks, (ii) head-level telemetry, and (iii) held-out language modeling perplexity. Our findings challenge the intuition that early induction circuit activation directly translates to better ICL. While Bi-Induct accelerates induction head emergence at smaller scales, this does not consistently yield better few-shot generalization. On standard LM benchmarks, Bi-Induct matches natural-only training; on function-style ICL probes, the 1B natural-only model performs best. Stress tests (e.g., label permutation, HITS@1 vs. HITS@3, 1 vs. 10 shots) preserve these trends. Telemetry reveals that larger models trained only on natural text develop broader and earlier-peaking induction heads, despite seeing no explicit induction patterns. Anti-induction data fails to elicit meaningful activation. Perplexity penalties from synthetic data shrink with scale, suggesting that larger models can absorb non-natural patterns with minimal cost. Crucially, ablating the top 2% of induction heads per layer degrades ICL more than random ablations, especially for natural-only models, indicating more centralized, load-bearing circuits. Bi-Induct variants exhibit more redundant induction activity, pointing to different circuit utilization patterns. Overall, we find that inducing activation is not sufficient: improvements in ICL hinge on whether these circuits become functionally necessary. These results underscore the importance of mechanism-aware pretraining diagnostics and data mixtures that foster load-bearing, not merely present, structure.
@@ 1 Introduction
Transformer language models learn a simple copy pattern early in training: when a token reappears in context, the model increases the probability of the token that followed the previous . Prior work identified a two-head motif implementing this behavior and linked it to in-context learning on pattern-matching tasks (Olsson et al., 2022). Despite its simplicity, this motif typically emerges only after many billions of tokens, well after the first training-loss plateau. Theoretical and empirical studies frame the delay as a phase transition (Chen et al., 2024; Edelman et al., 2024). Accelerating the onset of this transition could reduce compute and expose internal circuits for analysis earlier.
@@
A practical way to approach this is to intervene on the data rather than the architecture or objective. In contrast to synthetic-task-only plateau-shortening studies (Kim et al., 2025) and objective-level interventions such as multi-token prediction (Gloeckle et al., 2024), we adopt a data-rewrite perspective that is easy to deploy at scale: inject a small fraction of targeted inputs into the pretraining stream that selectively excite the putative induction mechanism while keeping the architecture and objective fixed. Concretely, we replace a small slice of natural tokens with synthetic copy snippets that cleanly exercise the copier-selector behavior of induction (forward copy) and anti-induction (backward copy). Copy-style cues are the canonical probe for the induction circuit and are widely used to measure it (Olsson et al., 2022; Nanda & Bloom, 2022). While other synthetic families have been studied (for example n-gram statistics (Edelman et al., 2024), p-hop tasks (Sanford et al., 2024), and intrinsic tasks (Gu et al., 2023)), copy snippets align most directly with the hypothesized mechanism and with distributional properties such as burstiness that correlate with the rise of in-context learning (Chan et al., 2022).
@@
These considerations lead to a single, testable question: Under iso-FLOPs, what is more effective for in-context learning—pure natural text, or natural text with a small fraction replaced by directional copy snippets that directly exercise the induction circuit?
@@
To answer this, we introduce Bi-Induct, a lightweight curriculum that interleaves short duplicate-span snippets with natural text during early training. We evaluate three variants under matched compute: forward induction, backward anti-induction, and a balanced mix where the direction is chosen independently at each injection. We assess the impact of our pretraining strategies on three axes: (i) downstream ICL performance on standard few-shot benchmarks; (ii) a mechanistic metric, namely the concentration of the top 2% most active attention heads per layer; and (iii) language modeling quality via held-out perplexity. We conduct experiments at three model scales (0.13B, 0.5B, 1B) under iso-FLOPs constraints, also using the 0.13B model as a design space to explore curriculum parameters, including data mix ratios, and span lengths effects.
Method and scope: We implement Bi-Induct as a lightweight curriculum, introducing a linearly annealed mixture of synthetic directional-copy data during early pretraining which we compare against a pure-natural baseline under iso-FLOPs conditions. To disentangle model capability from calibration, we evaluate few-shot ICL performance alongside two diagnostics: held-out perplexity (as a quality guardrail) and head-level telemetry measuring attention to copy patterns. Our objective is not to optimize benchmark scores, but to understand when and how targeted data augmentation induces load-bearing behavior for ICL.
@@
Matched-compute comparison across scales: We conduct an iso-FLOPs comparison of pure-natural pretraining vs. a small directional-copy replacement, run at 0.13B, 0.5B, and 1B parameters on the same corpus with identical token budgets and training steps per scale. We report per-task and composite few-shot ICL results alongside a perplexity guardrail and attention-head telemetry (forward/anti-copy scores) at the final checkpoint.
Key empirical finding: Early induction better ICL. While Bi-Induct accelerates and broadens induction-head activity at 0.13B and 0.5B parameters, this does not translate into superior ICL endpoints; at 1B, the natural-only baseline outperforms Bi-Induct on function-style probes.
Induction-head ablations revealing circuit load-bearing: Ablating the top 2% induction heads per layer reduces ICL more than matched random ablations at all scales. The largest relative drop is seen in the natural-only baseline, while Bi-Induct variants degrade less, suggesting a more redundant or distributed use of the induction circuit under synthetic augmentation.
Design-lab ablations (0.13B): We use the 0.13B model as a design lab to sweep curriculum parameters, including copy-span length, initial synthetic-to-natural ratio, and annealing schedules, to identify scaling-relevant settings.
ICL robustness diagnostics: We evaluate model robustness through shot-count variation, label- permutation stress tests, and decision-rule sensitivity via HITS@ on function-style probes.
@@ 2 Related Work
Induction heads and the mechanics of ICL: The induction-head motif—a two-head circuit that matches a repeated cue and copies its following token—was introduced by Olsson et al. (2022), who provided multiple converging tests linking it to the rise of in-context learning (ICL). Follow-up theory and controlled synthetic-task studies characterize the behavior as a phase transition: on Markov-chain data, models move from uniform predictions to unigram heuristics and then abruptly to bigram induction (Edelman et al., 2024). Provable analyses show that even shallow transformers implement generalized induction via a copier-selector-classifier pipeline, tightening the link between optimization dynamics and the circuit (Chen et al., 2024). At scale, targeted head ablations support causality: removing a small fraction of high-score induction heads reduces few-shot gains by up to 32% on abstract pattern tasks and weakens benefits on NLP tasks (Crosbie & Shutova, 2025). Open suites and tools (e.g., Pythia and TransformerLens) have made these effects reproducible across model sizes (Biderman et al., 2023; Nanda & Bloom, 2022).
Anti-induction and copy-suppression circuits: Beyond forward copying, models host heads that suppress copying or implement the backward, “anti-induction” direction. Work on negative heads explains copy suppression as a coherent mechanism that interacts with induction patterns (McDougall et al., 2023). Large-scale empirical reports find a pretrained asymmetry—transformers are stronger at forward induction than backward copy—an imbalance that targeted fine-tuning can reduce while isolating distinct head families (Veitsman et al., 2025). In parallel, Wang et al. (2025) mechanistically link the ‘repetition curse’ to over-dominant induction heads and propose head-level regularization to restore output diversity.
@@
Circuit discovery and emergence shaping: Mechanistic interpretability maps internal circuits via activation interventions (Zhang & Nanda, 2024), probing (Gurnee et al., 2023), and increasingly, sparse-autoencoder features (Cunningham et al., 2023). Our focus is earlier in the lifecycle: shaping the data distribution so that desirable circuits appear sooner and more predictably, then verifying the link with head-level telemetry.
Summary: Existing work investigates circuit discovery/emergence shaping and, separately, data rewriting, but rarely bridges the two—i.e., using mechanistic insight to design pretraining data and evaluating the outcome under matched-compute conditions. We make that link explicit: we target a canonical ICL circuit (forward and backward induction) with the minimal inputs that excite it (directional copy snippets), and compare pure natural pretraining to Bi-Induct, a small, linearly annealed replacement policy, under iso-FLOPs conditions on the same corpus. We measure effects behaviorally (few-shot ICL benchmarks) and mechanistically (top 2% head concentration by layer), alongside a perplexity guardrail. Unlike prior curricula that emphasize forward copy alone, we study a symmetric forward/backward curriculum side by side to ask whether targeted copy signals are more valuable than additional natural tokens at the same compute.
@@ 3 Bi-Induct
We investigate the effect of data rewrites on circuit emergence: we interleave synthetic copy snippets into the pretraining stream to explicitly teach the copy algorithm. Bi-Induct has two primary variants that differ only in the direction of the copy cue (forward vs. backward). A third variant, balanced, flips a coin between forward and backward injections to provide a mixed signal.
@@ 3.1 Synthetic snippet construction
Balanced/mix of forward and backward injections:
@@ 5.1 Downstream ICL Performance
Standard LM benchmarks: Figure 2(a) reports 5-shot macro-averages across 14 benchmarks. At each scale, at least one Bi-Induct variant matches or very slightly exceeds the natural-only baseline: at 0.13B, Anti is closest (22.5 vs. 22.7); at 0.5B, Induction leads (23.9 vs. 23.6); at 1B, Balanced is on par or marginally higher (24.3 vs. 24.2). Within measurement noise, copy-snippet curricula are largely performance-neutral at these scales (i.e., neither clearly degrading nor reliably improving downstream ICL).
@@ 5.2 Mechanistic Telemetry
(i) Emergence timing: At 0.13B and 0.5B, Bi-Induct variants show earlier induction-head emergence than the baseline (by roughly 3 and 2 layers, respectively). At 1B, the trend reverses: the baseline is the first to form clear induction peaks (around layers 10-11) and its early peaks are higher than any Bi-Induct variant. For anti-induction, absolute scores are small at all scales; the largest peak we observe is at 0.5B (Induction curriculum), followed by at 1B (Baseline). In keeping with Veitsman et al. (2025), forward-induction heads dominate in pretrained LMs; in our runs, even the Anti curriculum did not materially increase anti-induction copy scores. Notably, the strongest induction activity is concentrated in mid layers, in keeping with prior observations of where induction heads typically emerge (Olsson et al., 2022).
@@
Link to ICL performance:
These mechanistics results line up with the aggregate ICL results. On standard LM benchmarks (Figure 2(a)), macro ICL composites are similar across curricula at each scale: when Bi-Induct makes induction emerge earlier/stronger (0.13B/0.5B), performance is comparable to Baseline; when Baseline emerges earlier (1B), Bi-Induct is still comparable. A plausible explanation is a path shift: larger models can route more prediction mass through FFN/residual pathways, so head-peak magnitude and spread are not the sole determinants of few-shot accuracy on these tasks—especially because many of these benchmarks are knowledge and calibration-heavy, where FFN ‘key-value’ memories are known to contribute substantially (Geva et al., 2021). In contrast, on Todd et al. (2024)’s function-style suite (Figure 2(b)), the 1B Baseline’s earlier and broader induction spread coincides with a clear performance advantage over all Bi-Induct variants, suggesting these probes are more sensitive to the presence/strength of explicit copy heads.
Across scales, ablating the top-2% induction heads per layer at evaluation decreases the ICL composite more than ablating an equal number of random heads (Table 3). The effect is largest for the natural-only Baseline, and remains visible, though smaller, for all Bi-induct variants, a pattern that is consistent with either redundancy or a more distributed implementation of the behavior (random head ablations sometimes yield small improvements, as expected from noise/regularization effects). Detailed per-task comparisons for clean runs vs. induction-head and random-head ablations can be found in Table 12, Appendix E.
@@ 5.3 Guardrail: Language Modeling Perplexity
These observations validate our choice of a light, annealed Bi-Induct schedule: it preserves LM usability (competitive perplexity) and remains increasingly acceptable with scale, as the PPL gap consistently narrows across scale.

### v2

# Induction Signatures Are Not Enough: A Matched-Compute Study of Load-Bearing Structure in In-Context Learning

@@ Abstract
Mechanism-targeted synthetic data is increasingly proposed as a way to steer pretraining toward desirable capabilities, but it remains unclear how such interventions should be evaluated. We study this question for in-context learning (ICL) under matched compute (iso-FLOPs) using Bi-Induct, a lightweight data rewrite that interleaves short directional copy snippets into a natural pretraining stream: forward-copy (induction), backward-copy (anti-induction, as a directional control), or a balanced mix. Across 0.13B–1B decoder-only models, we evaluate (i) few-shot performance on standard LM benchmarks and function-style ICL probes, (ii) head-level copy telemetry, and (iii) held-out perplexity as a guardrail. Bi-Induct reliably increases induction-head activity, but this does not translate into consistent improvements in few-shot generalization: on standard LM benchmarks, Bi-Induct is largely performance-neutral relative to natural-only training, while on function-style probes the 1B natural-only model performs best. Despite explicit backward-copy cues, anti-induction scores remain near zero across scales, revealing a strong forward/backward asymmetry. Targeted ablations show a sharper distinction: removing the top 2% induction heads per layer harms ICL more than matched random ablations, with the largest relative drop occurring in the natural-only models. This indicates that natural-only training produces more centralized, load-bearing induction circuitry, whereas Bi-Induct tends to create more distributed and redundant induction activity. Our main conclusion is that eliciting a mechanism is not the same as making it load-bearing. For data-centric foundation model design, this suggests that synthetic data interventions should be evaluated not only by signature amplification, but by whether they create causally necessary computation while preserving natural-data modeling quality.
@@ 1 Introduction
Transformer language models learn a simple copy pattern early in training: when a token reappears in context, the model increases the probability of the token that followed the previous . Prior work identified a two-head motif implementing this behavior and linked it to in-context learning on pattern-matching tasks (Olsson et al., 2022). Despite its simplicity, this motif typically emerges only after many billions of tokens, well after the first training-loss plateau. Theoretical and empirical studies frame the delay as a phase transition (Chen et al., 2024; Edelman et al., 2024). In principle, if one could make induction-like computation become useful earlier in training, this could reduce the compute needed to reach ICL-relevant behaviors and would expose the responsible circuitry sooner for analysis.
@@
A practical way to approach this is to intervene on the data rather than the architecture or objective. In contrast to synthetic-task-only plateau-shortening studies (Kim et al., 2025) and objective-level interventions such as multi-token prediction (Gloeckle et al., 2024), we adopt a data-rewrite perspective that is easy to deploy at scale: inject a small fraction of targeted inputs into the pretraining stream that selectively excite the putative induction mechanism while keeping the architecture and objective fixed. Concretely, we replace a small slice of natural tokens with synthetic copy snippets that cleanly exercise the copier-selector behavior of induction (forward copy) and anti-induction (backward copy). Copy-style cues are the canonical probe for the induction circuit and are widely used to measure it (Olsson et al., 2022; Nanda and Bloom, 2022). While other synthetic families have been studied (for example n-gram statistics (Edelman et al., 2024), p-hop tasks (Sanford et al., 2024), and intrinsic tasks (Gu et al., 2023)), copy snippets align most directly with the hypothesized mechanism and with distributional properties such as burstiness that correlate with the rise of in-context learning (Chan et al., 2022).
@@
These considerations lead to a single testable question: under iso-FLOPs, is it more effective for ICL to pretrain purely on natural text, or to allocate a small early-training budget to synthetic directional copy snippets that directly exercise the induction circuit? Here, “earlier” refers to earlier-peaking induction signatures across layers at a fixed checkpoint under matched compute, not fewer optimization steps to reach a fixed capability.
@@
To answer this, we introduce Bi-Induct, a lightweight curriculum that interleaves short duplicate-span snippets with natural text during early training. We evaluate three variants under matched compute: forward induction, backward anti-induction, and a balanced mix where the direction is chosen independently at each injection. We assess each pretraining strategy on three axes: (i) downstream ICL performance on standard few-shot benchmarks and function-style probes; (ii) mechanistic telemetry targeting induction and anti-induction heads; and (iii) held-out perplexity as a quality guardrail. We study these effects across 0.13B, 0.5B, and 1B decoder-only models, using the 0.13B setting as a design lab for span length and mix-ratio selection before scaling the chosen operating point.
Our objective is not to present Bi-Induct as a generally superior pretraining recipe. Rather, we use it as a controlled case study for a broader data-centric question: when a synthetic data intervention successfully amplifies a target mechanism, does that mechanism become causally load-bearing for downstream behavior, or does it remain a redundant by-product of training? This distinction is especially important for data-centric foundation model design, where synthetic rewrites are attractive because they are easy to deploy, but their practical value depends on whether they improve useful computation rather than merely producing stronger internal signatures.
@@
A mechanism-aware evaluation criterion for synthetic data interventions. We distinguish between circuit emergence and circuit load-bearing: a targeted mechanism may become visible in telemetry without becoming necessary for task performance.
A matched-compute case study of this distinction. Under iso-FLOPs across 0.13B, 0.5B, and 1B models, we show that Bi-Induct reliably increases induction-head activity but does not reliably improve few-shot ICL, and at 1B the natural-only baseline performs best on function-style probes.
Causal evidence via targeted ablation. Using head-level copy telemetry together with top-2% induction-head ablations, we show that the largest ICL drops occur in the natural-only condition, indicating more centralized, load-bearing induction circuitry there than under Bi-Induct.
Directional controls for interpreting copy signals. By comparing forward induction, backward anti-induction, and a balanced mixture, we find a strong forward/backward asymmetry: even explicit anti-induction training does not materially increase anti-induction scores.
Practical lessons for the data-centric design of foundation models (FMs). Our results suggest that synthetic data interventions should be judged not only by whether they amplify a target signature, but by whether they create behaviorally useful and causally necessary computation without unduly sacrificing natural-data modeling quality.
@@ 2 Related Work
Induction heads and the mechanics of ICL: The induction-head motif-a two-head circuit that matches a repeated cue and copies its following token-was introduced by Olsson et al. (2022), who provided multiple converging tests linking it to the rise of in-context learning (ICL). Follow-up theory and controlled synthetic-task studies characterize the behavior as a phase transition: on Markov-chain data, models move from uniform predictions to unigram heuristics and then abruptly to bigram induction (Edelman et al., 2024). Provable analyses show that even shallow transformers implement generalized induction via a copier-selector-classifier pipeline, tightening the link between optimization dynamics and the circuit (Chen et al., 2024). At scale, targeted head ablations support causality: removing a small fraction of high-score induction heads reduces few-shot gains by up to 32% on abstract pattern tasks and weakens benefits on NLP tasks (Crosbie and Shutova, 2025). Open suites and tools (e.g., Pythia and TransformerLens) have made these effects reproducible across model sizes (Biderman et al., 2023; Nanda and Bloom, 2022).
Anti-induction and copy-suppression circuits: Beyond forward copying, models host heads that suppress copying or implement the backward, “anti-induction” direction. Work on negative heads explains copy suppression as a coherent mechanism that interacts with induction patterns (McDougall et al., 2023). Large-scale empirical reports find a pretrained asymmetry-transformers are stronger at forward induction than backward copy-an imbalance that targeted fine-tuning can reduce while isolating distinct head families (Veitsman et al., 2025). In parallel, Wang et al. (2025) mechanistically link the ‘repetition curse’ to over-dominant induction heads and propose head-level regularization to restore output diversity.
@@
Circuit discovery and emergence shaping: Mechanistic interpretability maps internal circuits via activation interventions (Zhang and Nanda, 2024), probing (Gurnee et al., 2023), and increasingly, sparse-autoencoder features (Cunningham et al., 2023). Our focus is earlier in the lifecycle: shaping the data distribution so that desirable circuits appear sooner and more predictably, then verifying the link with head-level telemetry.
Summary: Existing work investigates circuit discovery/emergence shaping and, separately, data rewriting, but rarely bridges the two-i.e., using mechanistic insight to design pretraining data and evaluating the outcome under matched-compute conditions. We make that link explicit: we target a canonical ICL circuit (forward and backward induction) with the minimal inputs that excite it (directional copy snippets), and compare pure natural pretraining to Bi-Induct, a small, linearly annealed replacement policy, under iso-FLOPs conditions on the same corpus. We measure effects behaviorally (few-shot ICL benchmarks) and mechanistically (top 2% head concentration by layer), alongside a perplexity guardrail. Unlike prior curricula that emphasize forward copy alone, we study a symmetric forward/backward curriculum side by side to ask whether targeted copy signals are more valuable than additional natural tokens at the same compute.
@@ 3 Bi-Induct
We investigate the effect of data rewrites on circuit emergence: we interleave synthetic copy snippets into the pretraining stream to explicitly exercise a canonical copy pattern associated with induction. Bi-Induct has two primary variants that differ only in the direction of the copy cue (forward vs. backward). A third variant, balanced, flips a coin between forward and backward injections to provide a mixed signal.
@@ 3.1 Synthetic snippet construction
Balanced (a mix of forward and backward injections):
@@ 5.1 Downstream ICL Performance
Standard LM benchmarks: Figure 2(a) reports 5-shot macro-averages across 14 benchmarks. At each scale, at least one Bi-Induct variant matches or very slightly exceeds the natural-only baseline: at 0.13B, Anti is closest (22.5 vs. 22.7); at 0.5B, Induction leads (23.9 vs. 23.6); at 1B, Balanced is on par or marginally higher (24.3 vs. 24.2). Within measurement noise, copy-snippet curricula are largely performance-neutral on standard LM benchmarks, but this neutrality is misleading in isolation: function-style probes and targeted ablations reveal important differences in whether induction becomes behaviorally load-bearing.
@@ 5.2 Mechanistic Telemetry
(i) Layerwise localization: At 0.13B and 0.5B, Bi-Induct variants show earlier induction-head emergence than the baseline (by roughly 3 and 2 layers, respectively). At 1B, the trend reverses: the baseline is the first to form clear induction peaks (around layers 10-11) and its early peaks are higher than any Bi-Induct variant. For anti-induction, absolute scores are small at all scales; the largest peak we observe is at 0.5B (Induction curriculum), followed by at 1B (Baseline). In keeping with Veitsman et al. (2025), forward-induction heads dominate in pretrained LMs; in our runs, even the Anti curriculum did not materially increase anti-induction copy scores. Notably, the strongest induction activity is concentrated in mid layers, in keeping with prior observations of where induction heads typically emerge (Olsson et al., 2022).
@@
Link to ICL performance (telemetry vs. causal reliance):
The layerwise copy telemetry above is correlational: it localizes where induction-like signatures concentrate, but it does not by itself establish that those heads are necessary for ICL. We therefore make a falsifiable distinction between emergence (telemetry) and load-bearing reliance (ablation sensitivity): if stronger/earlier signatures implied necessity, then removing the top induction heads would hurt most precisely in the conditions with the strongest telemetry peaks. Instead, we interpret telemetry jointly with behavioral results and the targeted ablations below.
On standard LM benchmarks (Figure 2(a)), macro ICL composites are similar across curricula at each scale. In particular, even when Bi-Induct shifts induction signatures toward earlier layers at 0.13B/0.5B, endpoints remain comparable to Baseline; and at 1B, despite the Baseline exhibiting an earlier depth-peak and a broader set of weakly-positive induction heads, Bi-Induct remains broadly competitive on these benchmarks. One plausible explanation is that many of these tasks are knowledge- and calibration-heavy, and larger models can route a substantial fraction of prediction mass through FFN/residual pathways rather than a small set of copy heads (consistent with evidence that FFNs act as key-value memories) (Geva et al., 2021). We treat this as a hypothesis rather than a mechanistic claim.
In contrast, on Todd et al. (2024)’s function-style suite (Figure 2(b)), the 1B Baseline shows a clear performance advantage, consistent with these probes being more sensitive to explicit copy-head computation. Notably, this advantage does not require that the Baseline have more uniformly high-scoring heads: visually (Figure 3), the Baseline is more concentrated (dominated by a small subset, including a single prominent head), whereas Bi-Induct variants can exhibit multiple prominent heads yet still fall short on these probes.
Causally, ablating the top-2% induction heads per layer decreases ICL composites more than ablating an equal number of random heads (Table 3). (Occasional small gains from random ablations are consistent with noise/regularization effects.) The relative drop is largest for the natural-only Baseline, while Bi-Induct variants degrade less, consistent with more redundancy or a more distributed implementation of induction-like behavior. A plausible explanation is front-loaded (in-training) recruitment: because Bi-Induct injects copy cues early and then anneals them away, it may recruit multiple induction-capable heads that persist as redundant “backup” pathways rather than becoming a single load-bearing bottleneck. Distinguishing early recruitment from late diffusion would require tracking copy telemetry across checkpoints. Detailed per-task comparisons for clean runs vs. induction-head and random-head ablations appear in Table 12, Appendix E.
@@ 5.3 Guardrail: Language Modeling Perplexity
These observations show that a light, annealed Bi-Induct schedule keeps the perplexity penalty bounded and shrinking with scale, but the natural-only baseline remains consistently better on held-out perplexity at every scale.
@@
5.4 Implications for Data-Centric Foundation Model Design
Taken together, these results suggest a practical evaluation principle for synthetic data interventions. Under matched compute, it is not sufficient to show that a data rewrite amplifies a target internal signature. A useful intervention should also improve or at least preserve downstream behavior, avoid unnecessary degradation in natural-data modeling quality, and ideally make the targeted computation more causally necessary rather than merely more visible. In our case, Bi-Induct succeeds at signature amplification but not at consistently creating load-bearing ICL circuitry, especially relative to natural-only training at 1B.